Solve the system: y^2=1-x^2
y^2=x^2-3x+2

Answer by mananth(16949)

y^2=1-x^2
y^2=x^2-3x+2
..
x^2-3x+2=1-x^2
2x^2-3x+1=0
2x^2-2x-x+1=0
2x(x-1)-1(x-1)=0
(x-1)(2x-1)=0
x=1 OR 1/2
plug the value of x to get value of y
y=0 OR sqrt 3/4